A rectangular plot of land has a length of 250m and a width of 175 m. The farmer wants to fence the plot using 6 strands of wire. What length of wire will s/he need?

Answer by checkley79(3341)

PERIMETER=2L+2W
P=2*250+2*175
P=500+350=850 FOR ONE STRAND OF WIRE.
850*6=5,100m FOR 6 STRANDS OF FENCING IS NEEDED.